OpenStudy (anonymous):

Solve, using linear combination. 3x + y = 4 2x + y = 4 (-1, 7) (0, 4) (-3, 12) (-3, 13)

OpenStudy (anonymous):

Well there are two options to use substitution method or elimination method, I'm sure you are aware of this.

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

I'll be using the substitution one in this case. 1) 3x + y = 4 I am going to rearrange to y =ax + b form y = -3x + 4 sub y = -3x + 4 into second equation 2x + y = 4 2x + (-3x + 4) = 4 2x -3x + 4 = 4 -x = 0 x = 0 sub x = 0 back into either equations to find y 2x + y = 4 2(0) + y = 4 0 + y = 4 y = 4 so (0,4) and the second answer.

OpenStudy (anonymous):

oooooo i get it now thank you :)

OpenStudy (anonymous):

no problem :) and remember you can always check with these type of questions if you have got it correct or not by subbing back the x and y values you obtained into the equations.

OpenStudy (anonymous):

ok thank you :)